-/5 POINTS MY NOTES Use Stokes theorem to evaluate | vxř. ñ ds where F = 9y?z, 6xz, 7x?y2 and S is the paraboloid z = x2 + ycalc hw- pls help!! (:
Answer #1

d+y=1 -2=1 Son 3 E = (942, 622, 7x²42) co x² + y²=1 ; 2=1 xi coso, dx=-sino y = sind, dy = coso A conding to stokes theorem S(4) F = 1644 2,34²2, 42) curl E li i i be by z Guyz 342 242 TER 13,0,0) plane: 2x+3y +2=6 - i (22-34²) -j (-bay) +Ê (-682) cu1 48 - loon +42²8 2² - -9 +332-1 an² + 3n3] + e [-x3-272+ art +27] de = [ 48* -50x?+4x!-2 **** (-4x+33 x-ıqd +7x4) +41-*_22x